if the roots of the equation ax^2 +bx +c=0 be (k+1)/k k+2/k+1 then (a+b+c)^2

Grade:11

1 Answers

Vikas TU
7 years ago
 (a+b+c)^2 = m
Take a^2 common then it would it become:
(1 – (-b/a) + (c/a) )^2 = m/a^2.
and also figure out as:
|x – y| = (b^2 – 4ac)/a^2 as one eqn. where x and y be roots.
write sum of roots and product of roots in termss of k and then try to solve.